# Few-Shot Joint Multimodal Entity-Relation Extraction via Knowledge-Enhanced Cross-modal Prompt Model

### v1

@@ Abstract.
Joint Multimodal Entity-Relation Extraction (JMERE) is a challenging task that aims to extract entities and their relations from text-image pairs in social media posts. Existing methods for JMERE require large amounts of labeled data. However, gathering and annotating fine-grained multimodal data for JMERE poses significant challenges. Initially, we construct diverse and comprehensive multimodal few-shot datasets fitted to the original data distribution. To address the insufficient information in the few-shot setting, we introduce the Knowledge-Enhanced Cross-modal Prompt Model (KECPM) for JMERE. This method can effectively address the problem of insufficient information in the few-shot setting by guiding a large language model to generate supplementary background knowledge. Our proposed method comprises two stages: (1) a knowledge ingestion stage that dynamically formulates prompts based on semantic similarity guide ChatGPT generating relevant knowledge and employs self-reflection to refine the knowledge; (2) a knowledge-enhanced language model stage that merges the auxiliary knowledge with the original input and utilizes a transformer-based model to align with JMERE’s required output format. We extensively evaluate our approach on a few-shot dataset derived from the JMERE dataset, demonstrating its superiority over strong baselines in terms of both micro and macro F1 scores. Additionally, we present qualitative analyses and case studies to elucidate the effectiveness of our model.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28–November 1, 2024; Melbourne, VIC, Australia.booktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28–November 1, 2024, Melbourne, VIC, Australiaisbn: 979-8-4007-0686-8/24/10doi: https://doi.org/10.1145/3664647.3680717ccs: Computing methodologies Information extractionccs: Computing methodologies Computer vision tasksccs: Information systems Users and interactive retrieval
@@ 1. Introduction
Recent focus has intensified on two pivotal subtasks for building a multimodal knowledge graph: Multimodal Named Entity Recognition (MNER)(Wu et al., 2020; Zheng et al., 2021c; Yu et al., 2020; Zhang et al., 2021b; Lu et al., 2022) and Multimodal Relation Extraction (MRE)(Zheng et al., 2021b, a; Zhao et al., 2023). These tasks aim to leverage user-generated content from social media, primarily composed of images and text, to generate structured knowledge. However, previous works treated these tasks independently and restricted the exploration of their interconnections (Wu et al., 2020; Zheng et al., 2021c; Yu et al., 2020; Zheng et al., 2021b, a; Zhao et al., 2023). To address this problem, Yuan et al. (2023) introduced a Joint Multimodal Entity-Relation Extraction (JMERE) task, aiming to enhance performance by integrating bidirectional interactions between the two subtasks. JMERE involves simultaneously extracting entities and their potential relations. However, previous JMERE work usually extracted features separately using pre-trained models (e.g., BERT for textes and ResNet for images), and employed multi-modal fusion layers for interaction, which typically require substantial training data (Yu and Zhang, 2022). Nevertheless, extensive multimodal data collection and annotation demand significant time and labor investments (Zhou et al., 2021). Furthermore, in real-world applications, access to labeled data is often constrained, and diverse domains necessitate specific datasets, presenting a challenge for data collection efforts. Thus, we first explore the JMERE task in a few-shot setting, aiming to advance the practical applications of multi-modal information extraction. To our knowledge, we are the first to focus on JMERE in a multimodal few-shot scenario (FS-JMERE).
@@
To address above challenges, we propose the Knowledge-Enhanced Cross-modal Prompt Model (KECPM) for Few-shot Joint Multimodal Entity-Relation Extraction (FS-JMERE), as shown in Fig. 3. Initially, to create a few-shot training and development dataset, we sampled data according to distinct relationship types’ data distribution. To address the insufficient information in FS-JMERE, we propose a method to acquire relevant knowledge from ChatGPT, which employs a dynamic prompt creation method and a knowledge selector to guide ChatGPT in generating more relevant auxiliary background knowledge. Formally, we manually annotate a limited set of human samples as candidate prompts, which comprises two primary components. The first part entails identifying the named entities and their relationships within the sentences. The second part involves offering comprehensive justifications by considering both the image and text content, along with relevant knowledge. Subsequently, KECPM employs a multimodal similarity-aware module to select human samples and integrates them into prompt templates fitted for the JMERE task, thereby introducing relevant background knowledge. To further improve the relevance and usefulness of knowledge generated by ChatGPT, we enable multiple rounds of self-reflection (Shinn et al., 2023), allowing it to rectify errors. Subsequently, we employ a knowledge selector to choose more accurate and relevant knowledge. This approach leverages ChatGPT’s few-shot learning capability and its capacity for self-reflection within in-context learning. The auxiliary knowledge from ChatGPT’s heuristic method is integrated with text and image captions, enhancing input for downstream models without requiring additional modality alignment for the FS-JMERE task. The main contributions of this study are summarized as follows:
@@ 2.2. Few-shot unimodal Information Extraction
Established few-shot learning approaches have been applied in unimodal information extraction (e.g., Named Entity Recognition (NER)(Fritzler et al., 2019; Wiseman and Stratos, 2019; Yang and Katiyar, 2020; Wang et al., 2022b) and Relation Extraction (RE) (Gao et al., 2019; Han et al., 2021; Liu et al., 2022a; Wu et al., 2023)). The works in Few-shot Information Extraction (FS-IE) can be divided into two categories: the first category uses limited data for supervised training. Fritzler et al. (2019) applied prototype networks (Snell et al., 2017) to address the few-shot NER (FS-NER) task. Inspired by nearest-neighbor inference in few-shot learning (Wiseman and Stratos, 2019), Yang et al. (Yang and Katiyar, 2020) employed supervised feature extractors for FS-NER. Within the few-shot RE (FS-RE) task, Gao et al. (2019) introduced attention mechanisms to augment prototype networks (Snell et al., 2017), and Han et al. (2021) incorporated relationship description information as a means of distinguishing between complex and straightforward tasks. Liu et al. (2022a) incorporate relationship description with support sample representations, resulting in notably superior results. Nevertheless, this approach fails to address the issue of insufficient information in the few-shot setting.
The second category focuses on pre-training or integrating external information to address the challenges of the few-shot setting. In the FS-NER task, Huang et al. (Huang et al., 2020) proposed a specialized noisy supervised pre-training approach, while Wang et al. (2021) explored model distillation techniques to enhance few-shot NER performance. In the FS-RE task, Qu et al. (2020) developed a comprehensive strategy involving the construction of a global relationship graph from Wikidata, which effectively facilitated the learning of posterior distributions of relation prototype vectors. Additionally, Yang et al. (2021) aimed to improve model performance by introducing entity concepts. However, constructing extensive unlabeled datasets and knowledge necessitates a significant human effort. Such an approach contradicts the initial goal of addressing the challenges posed by the few-shot setting. Besides, these methods have primarily been focused on the unimodal domain. Their direct application in the multi-modal domain has the risk of loss of valuable image-related information potentially.
@@ 4.6. Ablation Experiments
Finally, we attempted to incorporate visual object features (OF) from Mask R-CNN into the input of the language model, similar to previous work (Yang et al., 2023), to include more fine-grained information. However, this integration did not result in performance improvement. We consider that aligning features across different modalities from distinct pre-training datasets (text from T5, objects from Mask R-CNN) demands a significant number of data, which is inappropriate in the few-shot setting.
@@ 4.7. Analysis of Prompt Selection and Knowledge Reflection Modules
The results, as shown in Table 4, Firstly, the ChatGPT with few-shot settings has a performance gap with our proposed model. We have observed that ChatGPTgenerates entity and relationship types not present in the FS-JMERE pre-defined, such as ”building.” This underscores the importance of Stage 2 in our proposed method, as it effectively aligns with the FS-JMERE task while leveraging background knowledge from ChatGPT. Moreover, the PS module can notably enhance the quality of auxiliary knowledge generated by ChatGPT and improve the model performance. Besides, an optimal number of human-prompt instances can further refine the quality of auxiliary knowledge. However, merely incorporating human prompts without consistent improvement in performance may not suffice. Supplying more than 5 prompts (K=10) to ChatGPT may decrease the quality of auxiliary knowledge. In practice, it has been observed that an excessive number of manual examples introduces noise into the generation process of ChatGPT, potentially disrupting its original reasoning process and leading to the replication of manual examples.
@@ Acknowledgements.
Appendix A Effects of Different LLMs and LMMs
We evaluated large language models (Gemini 1.5, GLM4, GPT-4) and large multimodal models (GLM-4v, GPT-4o) on the FS-JMERE task, comparing them with the previous best results of ChatGPTk=5. The results are shown in Table 2.
Firstly, we observed that the performance of LMMs was significantly lower compared to LLMs. The best micro $F_{1}$ and macro $F_{1}$ scores for the LLMs were 25.34 and 19.04, respectively, whereas the multimodal models achieved 21.91 and 14.73, respectively. Possible reasons include: 1) LMMs have less textual knowledge compared to GPT-4 and GLM4, and 2) our prompts were designed for LLMs, so on the image side, we only input the test sample images without adding images from multiple manual prompts. As a result, the multimodal models might not have accurately captured the prompt intent.
Additionally, both GLM-4 and GPT-4 surpassed the original best ChatGPT performance when the number of prompts $K>5$. We incorporated the knowledge from GPT-4 with $K=10$ into our proposed method, KECPMGPT-4K=10, and achieved better performance. The scores of micro $F_{1}$ and macro $F_{1}$ increased from 39.96 and 21.29 to 42.24 and 24.33, respectively. GPT-4 also exhibited significantly fewer hallucination responses compared to ChatGPT, reducing them from 67 to 8 instances of meaningless responses such as ”There are no named entities or relations mentioned in the given text.” This improvement is likely due to GPT-4 and GLM-4 supporting up to 128K tokens, allowing for more comprehensive prompts and reducing hallucinations.
Model Micro Macro
P R F1 P R F1
Original Best ChatGPTK=5 21.81 19.01 20.31 14.44 16.29 15.30
Large Language Models Gemini-1.5K=1 25.20 21.30 23.08 12.75 13.11 12.92
Gemini-1.5K=5 24.90 24.20 24.54 14.70 16.20 15.44
Gemini-1.5K=10 26.55 23.10 24.7 14.52 17.06 15.67
GLM-4K=1 21.71 21.41 21.55 13.24 12.63 12.93
GLM-4K=5 24.34 23.94 24.14 15.54 17.71 16.55
GLM-4K=10 25.87 24.84 25.34 16.38 18.35 17.31
GPT-4K=1 17.80 18.32 18.05 11.88 10.31 11.04
GPT-4K=5 21.12 21.42 21.26 19.84 16.40 17.95
GPT-4K=10 24.78 24.42 24.60 20.60 17.70 19.04
Multimodal Large Models GLM-4vK=1 16.21 12.00 13.79 9.89 7.49 8.52
GLM-4vK=5 18.71 17.52 18.13 10.32 7.79 8.88
GLM-4vK=10 18.07 4.87 7.67 8.03 4.24 5.54
GPT-4oK=1 19.32 15.47 17.18 12.18 9.42 10.62
GPT-4oK=5 20.15 0.20 20.07 12.69 9.87 11.10
GPT-4oK=10 22.15 21.68 21.91 13.44 16.29 14.73
KECPM (our method) ChatGPTK=5 44.06 36.55 39.96 24.45 18.85 21.29
GPT-4K=10 46.51 38.69 42.24 28.44 21.26 24.33
Table 5. Results of Different LLMs and LMMs in the FS-JMERE Task.
Appendix B Prompt template

### v2

@@ Abstract.
Joint Multimodal Entity-Relation Extraction (JMERE) is a challenging task that aims to extract entities and their relations from text-image pairs in social media posts. Existing methods for JMERE require large amounts of labeled data. However, gathering and annotating fine-grained multimodal data for JMERE poses significant challenges. Initially, we construct diverse and comprehensive multimodal few-shot datasets fitted to the original data distribution. To address the insufficient information in the few-shot setting, we introduce the Knowledge-Enhanced Cross-modal Prompt Model (KECPM) for JMERE. This method can effectively address the problem of insufficient information in the few-shot setting by guiding a large language model to generate supplementary background knowledge. Our proposed method comprises two stages: (1) a knowledge ingestion stage that dynamically formulates prompts based on semantic similarity guide ChatGPT generating relevant knowledge and employs self-reflection to refine the knowledge; (2) a knowledge-enhanced language model stage that merges the auxiliary knowledge with the original input and utilizes a transformer-based model to align with JMERE’s required output format. We extensively evaluate our approach on a few-shot dataset derived from the JMERE dataset, demonstrating its superiority over strong baselines in terms of both micro and macro F1 scores. Additionally, we present qualitative analyses and case studies to elucidate the effectiveness of our model. Code and Data are released at https://github.com/YuanLi95/KECPM.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28–November 1, 2024; Melbourne, VIC, Australia.booktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28–November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3680717isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Information extractionccs: Computing methodologies Computer vision tasksccs: Information systems Users and interactive retrieval
@@ 1. Introduction
Recent focus has intensified on two pivotal subtasks for building a multimodal knowledge graph: Multimodal Named Entity Recognition (MNER)(Wu et al., 2020; Zheng et al., 2021c; Yu et al., 2020; Zhang et al., 2021b; Lu et al., 2022) and Multimodal Relation Extraction (MRE)(Zheng et al., 2021b, a; Zhao et al., 2023). These tasks aim to leverage user-generated social media content, primarily images and text, to generate structured knowledge. However, previous works treated these tasks independently and restricted the exploration of their interconnections (Wu et al., 2020; Zheng et al., 2021c; Yu et al., 2020; Zheng et al., 2021b, a; Zhao et al., 2023). To address this problem, Yuan et al. (2023) introduced a Joint Multimodal Entity-Relation Extraction (JMERE) task, aiming to enhance performance by integrating bidirectional interactions between the two subtasks. JMERE involves simultaneously extracting entities and their potential relations. However, previous JMERE work usually extracted features separately using pre-trained models (e.g., BERT for textes and ResNet for images), and employed multi-modal fusion layers for interaction, which typically require substantial training data (Yu and Zhang, 2022). Nevertheless, extensive multimodal data collection and annotation demand significant time and labor investments (Zhou et al., 2021). Furthermore, in real-world applications, access to labeled data is often constrained, and diverse domains necessitate specific datasets, presenting a challenge for data collection efforts. Thus, we first explore the JMERE task in a few-shot setting, aiming to advance the practical applications of multi-modal information extraction. To our knowledge, we are the first to focus on JMERE in a multimodal few-shot scenario (FS-JMERE).
@@
To address above challenges, we propose the Knowledge-Enhanced Cross-modal Prompt Model (KECPM) for Few-shot Joint Multimodal Entity-Relation Extraction (FS-JMERE), as shown in Fig. 3. Initially, to create a few-shot training and development dataset, we sampled data according to distinct relationship types’ data distribution. To address the insufficient information in FS-JMERE, we propose a method to acquire relevant knowledge from ChatGPT, which employs a dynamic prompt creation method and a knowledge selector to guide ChatGPT in generating more relevant auxiliary background knowledge. Formally, we manually annotate a limited set of human samples as candidate prompts, which comprises two primary components. The first part entails identifying the named entities and their relationships within the sentences. The second part involves offering comprehensive justifications by considering both the image and text content, along with relevant knowledge. Subsequently, KECPM employs a multimodal similarity-aware module to select human samples and integrates them into prompt templates fitted for the JMERE task, thereby introducing relevant background knowledge. To further improve the relevance and usefulness of knowledge generated by ChatGPT, we enable multiple rounds of self-reflection (Shinn et al., 2023; Yuan et al., 2024), allowing it to rectify errors. Subsequently, we employ a knowledge selector to choose more accurate and relevant knowledge. This approach leverages ChatGPT’s few-shot learning capability and its capacity for self-reflection within in-context learning. The auxiliary knowledge from ChatGPT’s heuristic method is integrated with text and image captions, enhancing input for downstream models without requiring additional modality alignment for the FS-JMERE task. The main contributions of this study are summarized as follows:
@@ 2.2. Few-shot unimodal Information Extraction
Established few-shot learning approaches have been applied in unimodal information extraction (e.g., Named Entity Recognition (NER)(Fritzler et al., 2019; Wiseman and Stratos, 2019; Yang and Katiyar, 2020; Wang et al., 2022b) and Relation Extraction (RE) (Gao et al., 2019; Ren et al., 2020; Han et al., 2021; Liu et al., 2022a; Wu et al., 2023)). The works in Few-shot Information Extraction (FS-IE) can be divided into two categories: the first category uses limited data for supervised training. Fritzler et al. (2019) applied prototype networks (Snell et al., 2017) to address the few-shot NER (FS-NER) task. Inspired by nearest-neighbor inference in few-shot learning (Wiseman and Stratos, 2019), Yang et al. (Yang and Katiyar, 2020) employed supervised feature extractors for FS-NER. Within the few-shot RE (FS-RE) task, Gao et al. (2019) introduced attention mechanisms to augment prototype networks (Snell et al., 2017), and Han et al. (2021) incorporated relationship description information as a means of distinguishing between complex and straightforward tasks. Liu et al. (2022a) incorporate relationship description with support sample representations, resulting in notably superior results. Nevertheless, this approach fails to address the issue of insufficient information in the few-shot setting.
The second category focuses on pre-training or integrating external information to address the challenges of the few-shot setting. In the FS-NER task, Huang et al. (Huang et al., 2020) proposed a specialized noisy supervised pre-training approach, while Wang et al. (2021) explored model distillation techniques to enhance few-shot NER performance. In the FS-RE task, Qu et al. (2020) developed a comprehensive strategy involving the construction of a global relationship graph from Wikidata, effectively facilitating the learning of posterior distributions of relation prototype vectors. Additionally, Yang et al. (2021) aimed to improve model performance by introducing entity concepts. However, constructing extensive unlabeled datasets and knowledge necessitates a significant human effort. Such an approach contradicts the initial goal of addressing the challenges posed by the few-shot setting. Besides, these methods have primarily been focused on the unimodal domain. Their direct application in the multi-modal domain has the risk of loss of valuable image-related information potentially.
@@ 4.6. Ablation Experiments
Finally, we attempted to incorporate visual object features (OF) from Mask R-CNN into the input of the language model, similar to previous work (Yang et al., 2023), to include more fine-grained information. However, this integration did not result in performance improvement. We consider that aligning features across different modalities from distinct pre-training datasets (text from T5, objects from Mask R-CNN) demands significant data, which is inappropriate in the few-shot setting.
@@ 4.7. Analysis of Prompt Selection and Knowledge Reflection Modules
As shown in Table 4, the ChatGPT with few-shot settings has a performance gap with our proposed model. We have observed that ChatGPTgenerates entity and relationship types not present in the FS-JMERE pre-defined, such as ”building.” This underscores the importance of Stage 2 in our proposed method, as it effectively aligns with the FS-JMERE task while leveraging background knowledge from ChatGPT. Moreover, the PS module can notably enhance the quality of auxiliary knowledge generated by ChatGPT and improve the model performance. Besides, an optimal number of human-prompt instances can further refine the quality of auxiliary knowledge. However, merely incorporating human prompts without consistent improvement in performance may not suffice. Supplying more than 5 prompts (K=10) to ChatGPT may decrease the quality of auxiliary knowledge. In practice, it has been observed that an excessive number of manual examples introduces noise into the generation process of ChatGPT, potentially disrupting its original reasoning process and leading to the replication of manual examples.
@@ Acknowledgements.
Appendix A Prompt Template